2) Luke is on the Track and Field team and Angeline is on the weightlifting team at their community college. Luke can run a mile in 248 seconds, the mean mile time on the team is 273 seconds with a standard deviation of 15 seconds. Angeline can lift (dead lift) 335 lbs, whereas the team can lift a mean of 315 lbs and varies with a standard deviation of 9.9 lbs. Luke and Angeline are trying to decide who is a better athlete when compared to the other athletes on their team. In weightlifting the goal is to lift more, but when a running a mile you want a smaller/faster time. Decide who the better athlete is and support with statistical analysis (hint: use z-scores).
Answer #1

Answer:

z = (x-µ)/σ

z value for Luke is z = (248-273)/15 = -1.67

z value for Angeline is z = (335-315)/9.9 = 2.02

Angeline z score 2.02 is far from mean 0 value than Luke z value -1.67.

Angeline is better athlete than Luke.